### v1

# Performance Assessment of Universal Machine Learning Interatomic Potentials: Challenges and Directions for Materials’ Surfaces

Abstract
Machine learning interatomic potentials (MLIPs) are one of the main techniques in the materials science toolbox, able to bridge ab initio accuracy with the computational efficiency of classical force fields. This allows simulations ranging from atoms, molecules, and biosystems, to solid and bulk materials, surfaces, nanomaterials, and their interfaces and complex interactions. A recent class of advanced MLIPs, which use equivariant representations and deep graph neural networks, is known as universal models. These models are proposed as foundational models suitable for any system, covering most elements from the periodic table. Current universal MLIPs (UIPs) have been trained with the largest consistent dataset available nowadays. However, these are composed mostly of bulk materials’ DFT calculations. In this article, we assess the universality of all openly available UIPs, namely MACE, CHGNet, and M3GNet, in a representative task of generalization: calculation of surface energies. We find that the out-of-the-box foundational models have significant shortcomings in this task, with errors correlated to the total energy of surface simulations, having an out-of-domain distance from the training dataset. Our results show that while UIPs are an efficient starting point for fine-tuning specialized models, we envision the potential of increasing the coverage of the materials space towards universal training datasets for MLIPs.
I Introduction
Machine learning (ML) and artificial intelligence (AI) are rapidly changing how many aspects of our society evolve. In particular, these new approaches and techniques are added to the scientists’ toolbox, and combined with traditional ones, are increasingly demonstrating novel ways to advance scientific challenges. One of the most impressive advances in recent materials science has been the use of artificial intelligence to improve and accelerate the process of materials discovery, design, and understanding [1, 2, 3, 4, 5]. The field of machine learning interatomic potentials (MLIPs) has benefited the most from these advances, being currently one of the largest and most mature areas successfully using machine learning [6, 7, 8, 9].
MLIPs enable us to bridge the gap of computational simulations with both ab initio accuracy (typically at the DFT level of theory) and traditional force-field computational efficiency. They are capable of succeeding by going beyond traditional few-parameter analytical forms for describing the potential energy surfaces (PES) of systems. This is done by combining flexible machine-learned functional forms and algorithms with features or descriptors that are particularly adequate for these data-driven methods. To train such flexible and higher complexity functions, high-quality and high-volume datasets are mandatory [10], with datasets sufficiently representing the domain of interest. Even greater care is needed for generalization not only to larger simulations but also to a wide range of different systems than those of the reference dataset [11], known as extensibility and transferability, respectively. [9]
Recently, a class of MLIPs employing deep graph networks [12, 13, 14] has improved previous results in a range of systems and benchmarks. With the use of E(3)-equivariant representations, models have achieved not only accuracy but also unprecedented data efficiency, such as the Neural Equivariant Interatomic Potentials (NequIP) [15]. These advances led the community to envision universal MLIPs (UIPs), general models able to simulate any system across most of the periodic table [16]. These universal potentials have been proposed as foundational models, a starting point from which further specialized models could be trained upon.
Chronologically, these models improved by growing both their datasets and number of parameters: M3GNet [17] was trained with 188k structures and has 228k params; CHGNet [18] uses atomic magnetic moments as additional representation inputs and introduced the MPtrj dataset with 1.58M structures from the Materials Project database and has 413k params; MACE MP-0 [19, 20, 21] was also trained with MPtrj and has 4.69M params; and the proprietary GNoME model [22] was trained with an active learning dataset of 89M structures starting from the Materials Project, and has 16.2M parameters. As usual in the deep learning community, a catalyst for advances is the creation of benchmarks to assess performance in a comparable task, the Matbench Discovery [23] serves this purpose in the task of materials discovery for universal models. However, the models are trained on the largest systematic DFT databases available to date, which include millions of bulk materials simulations but are currently not comprehensible in the complete materials space of all possible combinations of atoms, compositions, and structures, such as molecules, surfaces and interfaces, nanomaterials, and so on.
Here we investigate one of the main questions for users of MLIPs: given a new problem, what is the most accurate and efficient way to tackle it? The assessment follows two steps, i) evaluating if universal MLIPs are accurate enough for systems that are significantly different than the training dataset, with the representative example case of surfaces, and ii) if it is better to train a specialized MLIP from scratch or to fine-tune a universal foundational model. We show that currently available UIPs do not display sufficient zero-shot performance in this representative task. In terms of efficiency, fine-tuning UIPs can accelerate training and accuracy by incorporating alchemical transferable knowledge, thus requiring only a modest dataset to achieve sufficient accuracy in specialized tasks.
II Results and Discussion
II.1 Dataset exploration
We start by querying the Materials Project (MP) database for data on surfaces. This data is available for unary systems [24]. We gathered a total of 1497 different surface structures that are generated from a total of 138 different bulk systems, comprehending 73 different chemical elements.
Figure 1: MP surfaces dataset exploration. (a) Periodic table heatmap for the number of surfaces of each element. (b) Boxplot and violin plot for surface energy of each element within the dataset. The horizontal lines in the middle of the boxes mark the medians. The boxes are plotted from the first to the third quartile. The whiskers extend to $1.5$ times the interquartile range from the third (above) and first (below) quartile and empty circles mark the outliers (data points outside the whiskers).
Figure 1 shows the distribution of surface data available on Materials Project for each element. Most materials are contemplated, however, noble gases, rare-earth elements, and also heavy synthetic elements are not present, see Fig. 1a. The alkali metals and alkaline earth metals are the groups with the most representation, followed by carbon materials. From Fig. 1b, we observe that the median surface energy increases from the edges to the center of the same period in the periodic table. The same is not observed for rare-earth elements (see the Supplemental Information). Overall the spread of surface energies is small, except for carbon materials, which also show significant outliers.
II.2 Inference of universal machine learning atomistic potentials
We proceed with an assessment of the universal MLIPs (UIPs) in their ability to predict the surface energy data of the materials under consideration. Initially, we gauge their performance in predicting the total energy of the bulk structures. Given their nature as universal MLIPs trained on bulk data sourced from the Materials Project, a high level of accuracy is anticipated for these specific structures.
[…]
Figure 2b illustrates the prediction error for the total energy of surfaces. As these structures differ from the models’ training set, the error rate increases. Surprisingly, CHGNet exhibits greater accuracy than both MACE and M3GNet in this context, with M3GNet also surpassing MACE. This observation highlights the inherent trade-off between accuracy and generalization. Although MACE performs better on the test set, its predictive capability diminishes when extrapolating to unknown data dissimilar to the training set.
Figure 2: Performance assessment of the universal interatomic potentials over the surfaces dataset. (a) Parity plot for the total energy per atom of the bulk systems that gave origin to the surfaces of the dataset. (b) Parity plot for the total energy per atom of the surfaces within the dataset. (c) Parity plot for the surface energy ($\gamma_{hkl}^{\sigma}$) for the surfaces within the dataset. For (a), (b), and (c), the dashed line marks the $x=y$ line. (d) Boxplot and violin plot for the error in the prediction of the surface energy ($\gamma_{hkl}^{\sigma}$) from the three universal interatomic potentials evaluated. The horizontal lines in the middle of the boxes mark the medians. The boxes are plotted from the first to the third quartile. The whiskers extend to $1.5$ times the interquartile range from the third (above) and first (below) quartile and empty circles mark the outliers (data points outside the whiskers).
Next, we evaluate these UIPs on predicting the surface energy. The surface energy $\gamma_{hkl}^{\sigma}$ of a surface is defined by its Miller indexes $(\mathrm{h\,k\,l})$ and termination $\sigma$ as calculated from the following expression
$\gamma_{hkl}^{\sigma}=\frac{E_{\rm slab}^{hkl,\sigma}-n_{\rm slab}^{hkl,\sigma}\epsilon_{\rm bulk}}{2A_{\rm slab}^{hkl,\sigma}}$ (1)
where $E_{\rm slab}^{hkl,\sigma}$ is the surface (slab) total energy, $n_{\rm slab}^{hkl,\sigma}$ is the number of sites in the surface slab, $\epsilon_{\rm bulk}$ is the bulk total energy per atom, and $A_{\rm slab}^{hkl,\sigma}$ is the area of the surface defined by the $(\mathrm{h\,k\,l})$ indexes.
Figure 2c shows the parity plot for these models and Fig. 2d their error distribution. In the case of surface energy, CHGNet is still the most accurate, however MACE follows in second and M3GNet in third. CHGNet mostly underestimates the surface energy while MACE shows several cases of overestimation. Like CHGNet, M3GNet underestimates the surface energy. Ideally, accuracy over bulk and surface total energy is desired to achieve accurate surface energies, however, error cancellation also plays a part in this analysis, therefore keeping in comparison the bulk and surface total energy is essential to gain insight into these models’ performance. Additionally, we do not observe any correlation between the predictions of each model (see the Supplemental Information), that is, the errors are inherent from each model’s limits due to their different training set, representations, and architecture.
II.3 Understanding errors
II.3.1 System chemistry
Figure 3: Root mean squared error (RMSE) of the universal interatomic potentials concerning surface chemistry. The upper triangle is surface energy ($\gamma_{hkl}^{\sigma}$) RMSE, the lower leftmost triangle is the bulk total energy per atom RMSE and the lower rightmost triangle is the surface total energy per atom RMSE. The numbers in parenthesis are the number of surface structures evaluated for each chemical element.
Next, we show the RMSE of these models separated by system chemistry. Figure 3 shows a comparison between the RMSE on the surface energy and also bulk and surface total energy per atom, which we take as the limiting cases for the accuracy of each model. Figure 3 illustrates that errors in bulk predictions are not consistently correlated with errors in surface energy predictions. For example, while MACE shows poor performance in predicting the surface energy of tungsten (W), its bulk predictions are relatively accurate. Conversely, for caesium (Cs), both bulk and surface total energy predictions are accurate, but the surface energy predictions are not, which is a result of error cancellation. Despite these discrepancies, MACE demonstrates small errors for several other systems, for instance, copper (Cu) is well represented by MACE. Likewise, M3GNet exhibits a higher frequency of large errors, particularly for elements such as boron (B), carbon (C), manganese (Mn), iron (Fe), niobium (Nb), molybdenum (Mo), technetium (Tc), ruthenium (Ru), rhenium (Re), osmium (Os), iridium (Ir), and platinum (Pt). CHGNet shows fewer problematic predictions compared to M3GNet. Moreover, elements such as carbon (C), manganese (Mn), tungsten (W), iron (Fe), rhenium (Re), osmium (Os), and iridium (Ir) also require careful consideration, as these elements also present challenges for the UIPs. The comparison with the surface total energy shows that they are the major root for the deviations of the surface energy predictions, as expected. Surfaces are not included in the training set of these universal models, therefore they struggle with extrapolating from the bulk structures.
II.3.2 System global geometries (surfaces structures)
Figure 4: Root mean squared error (RMSE) of the universal interatomic potentials concerning surface orientation $(h\,k\,l)$. The numbers in parenthesis are the number of surface structures evaluated for each Miller index.
Figure 4 shows the RMSE of each UIP separated by surface $(h\,k\,l)$ index. Although one expects that lower-angle surfaces show fewer errors because the atoms would keep a higher coordination degree, this depends on the termination of the slab, therefore extracting correlations with the surface orientation in such case is challenging. Across the different surface orientations, all UIPs show slightly similar errors in their prediction. Several cases of hexagonal systems appear to have a higher error than cubic ones. However, higher errors are spread across the different orientations.
II.3.3 System local geometries (local environments, beyond surfaces structures)
Figure 5 shows a kernel PCA (kPCA) map created to represent each structure and its chemistry accounting for their local environments. The descriptor used for each structure is created using MACE’s representation with the same hyperparameters used for training the universal model, that is, a radius cutoff of $6.0\;\text{\AA}$ and three layers of 128 channels with a maximum angular momentum of $l_{\rm max}=2$. For each structure, we use a simple average between the atomic sites, and the comparison between structures is done by the REMatch kernel [26].
Figure 5 shows the kPCA map colored by the absolute error given by the MACE model for surface energy. Interestingly, the MACE representation is enough to single out most systems. The representation seems to take great advantage of the different chemical species within the dataset. As shown in Fig. 3, tungsten (W) presents most of the error from MACE. Additionally, in Fig. 5 we also try to associate the kPCA map with the lattice of each original bulk structure. We highlight the surfaces that originate from face-centered cubic (FCC), body-centered cubic (BCC), and hexagonal closed-packed (HCP) lattices. The error is not directly dependent on the lattice symmetry. Detailed analysis of each system chemistry on the kPCA map is available in the Supporting Information.
Figure 5: Kernel PCA (kPCA) map of the surfaces dataset, where each point corresponds to a structure and distances represent differences in input features. kPCA map with MACE representation and REMatch kernel, colored by the absolute surface energy error given by the MACE model, with selected elements highlighted. Markers represent the lattice structure of the bulk that generated each surface, with FCC, BCC, and HCP crystalline systems highlighted.
II.4 Comparison: universal vs specialized MLIPs
II.4.1 Universal vs NequIP/MTP
Our next step in this assessment is to address the question: “If one is interested in materials’ surfaces, are universal interatomic potentials a good starting point, or should one train a specialized interatomic potential for the task at hand?”. For this comparison, we started by selecting candidates for specialized MLIPs. Keeping in mind the performance-computational cost trade-off, the moment tensor potential (MTP) [27] formalism is ideal since it can attain very low test errors with a rather small dataset and number of parameters [25, 28]. Zuo et al. [25] performed a systematic benchmark of “pre-universal” MLIPs and made available not only the training and test sets of their performance and computational cost assessment, but also the models they have trained for the elements: nickel (Ni), copper (Cu), lithium (Li), molybdenum (Mo), silicon (Si), and germanium (Ge). We called MTP [JPCA] the set of MTP models as available from ref. [25] for each of the elements considered in their work. We trained another set of MTP models for these elements, however, using the bulk structures as queried from the Materials Project database. We called this set of models MTP [$*{}^{*}$], as this is derived from a reduced version of the MPtrj dataset, that is, for these selected elements, as in ref. [25]. Mind that for both MTP [JPCA] and MTP [$*{}^{*}$], there is a model for each element, while for UIPs, a single model handles all the elements.
We added the Neural Equivariant Interatomic Potential (NequIP) to our comparison as they are equivariant neural networks, similar to the UIPs, and highly efficient in data use. In this case, we trained two sets of NequIP models, one using ref. [25] training and test sets, which we called NequIP [JPCA], and a second set of models, as we did with MTP, using the bulk structures as queried from the Materials Project for the same elements used in ref. [25], which we call NequIP [$*{}^{*}$]. Our final comparison contrasts the UIPs: MACE, M3GNet, and CHGNet with four sets of specialized MLIPs: MTP [JPCA], MTP [$*{}^{*}$], NequIP [JPCA], and NequIP [$*{}^{*}$], as shown in Fig. 6 and summarized in Table 1. See the Supporting Information for the parity plot containing all these models for this reduced set of elements.
First, by comparing the RMSE in Table 1 with the RMSE of the full set of elements from the MPtrj dataset, the RMSE of UIPs shows a different trend. For this smaller set of elements, although MACE is still not the most accurate total energy prediction on the bulk structures, there is not a significant increase in the error for surface total energy prediction. M3GNet on the other hand shows a significant error increase. CHGNet however follows the opposite trend, with surface total energy errors being smaller than bulk total energy errors. For surface energy, MACE is the most accurate between the UIPs.
Comparing first MTP and NequIP with the two different datasets: JPCA and $*{}^{*}$. The models trained with the JPCA dataset present an overall better performance. We attribute this to be sheer influence of the construction of the dataset. The JPCA dataset is more diverse in types of structures containing: MD trajectory snapshots, defects, deformations, and even a few surfaces, while the $*{}^{*}$ datasets contain only bulk structures at their relaxed geometries and relaxation trajectories. Comparison between MTP and NequIP for the same dataset brings the evident power of MTP. Although NequIP showcases smaller RMSE, MTP is not far behind, especially for bulk structures. However, as we go to surfaces, NequIP showcases higher generalization power. Comparison between specialized and universal IPs demonstrates that UIPs lack the generalization power to yield reasonable predictions for surfaces. This is the case when taking into consideration both MTP and NequIP models trained with the JPCA dataset. As expected, the lack of diversity on the $*{}^{*}$ dataset hinders the performance of specialized MLIPs.
Figure 6: Performance comparison between universal, specialized, and universal fine-tuned interatomic potentials. The root mean squared errors (RMSE) of the universal interatomic potentials are compared in relation to the surface chemistry, for selected elements. The upper triangle is the surface energy ($\gamma_{hkl}^{\sigma}$) RMSE, the lower leftmost triangle is the bulk total energy per atom RMSE and the lower rightmost triangle is the surface total energy per atom RMSE. The numbers in parenthesis are the number of surface structures evaluated for each chemical element.
Table 1: Root mean squared errors (RMSE) for universal interatomic potentials, specialized machine learning interatomic potentials, and fine-tuned universal interatomic potential for a selected set of elements: Ni, Cu, Li, Mo, Si, and Ge. For the column type, we label the models as universal interatomic potential (UIP), local representation interatomic potential (LRIP), graph neural network (GNN), and UIP fine-tuned (UIP-FT).
RMSE over structures with Ni, Cu, Li, Mo, Si, and Ge
ML model
Bulk total energy
[…]
Surface total energy
[…]
Surface energy
[…]
Type
MACE 0.0471 0.0488 0.3974 UIP
M3GNet 0.0227 0.0979 0.6920 UIP
CHGNet 0.0791 0.0386 0.5128 UIP
MTP [JPCA] 0.0369 0.0480 0.2715 LRIP
MTP [$*{}^{*}$] 0.0519 0.4281 2.7932 LRIP
NequIP [JPCA] 0.0372 0.0181 0.2097 GNN
NequIP [$*{}^{*}$] 0.0316 0.0762 0.6511 GNN
MACE [FT] 0.0367 0.0496 0.2183 UIP-FT
As demonstrated in Fig. 6, for most elements where UIPs show an increased error, MTP [JPCA] and NequIP [JPCA] demonstrate high-accuracy predictions. Predictions for systems such as Mo and Si show great improvement from the specialized MLIPs. Systems containing Ge on the other hand show worse performance for bulk structures while slightly improved performance for surface structures.
II.4.2 Universal (zero-shot) vs Universal (fine-tuned)
Since UIPs seem to not be great at dealing with surfaces, we now address the question: “What can we do to improve them?”. One route to improving general models such as the UIPs is fine-tuning, which consists of continuing the training of these models however with a different dataset, tailored to make these universal models more specialized in particular tasks. In our case, we can use the training and test sets of ref. [25] to fine-tune these UIPs. By introducing a more diverse set of structures for each chemical element, we expect that the UIP will improve in predicting surface (total) energies while taking advantage of what it has already learned from the MPtrj dataset. We choose to fine-tune the MACE model, given the availability and ease of use of the corresponding implementation. We fine-tune our MACE model for another 200 epochs using the training set from ref. [25]. The fine-tuned model is called MACE [FT], as shown in the last line of Fig. 6 and Table 1.
After fine-tuning, we observe that the prediction of bulk total energies largely improved for most systems. Is noticeable how fine-tuning should be performed with care. As discussed before, both MTP and NequIP did not show great performance over bulk structures composed of germanium (Ge). After fine-tuning, MACE [FT] decreased its performance on germanium (Ge) bulk structures. The performance on surface (total) energies improved for most elements. On nickel (Ni) the RMSE slightly increased, and for copper (Cu) the increase was more significant. Overall, the fine-tuned universal model represents an important improvement: for the selected systems, the fine-tuned MACE model is of similar performance to specialized MLIPs.
III Conclusions
We have assessed one of the main questions for working with MLIPs: given a new problem, what is the best and most efficient way to tackle it? The answer follows at least two steps, evaluating if MLIPs are accurate enough and if not, if it is better to train a specialized MLIP from scratch or to fine-tune a universal foundational model. From our results for surface energies, we see that the total energies for surface geometries are modestly accurate, however, not good enough for specific properties. Therefore, as expected, the most important aspect is the domain distance of the training dataset to the target space, composed of both chemistry and structure.
In terms of efficiency, fine-tuning UIPs can greatly accelerate the training by incorporating alchemical transferable knowledge from the large bulk dataset (in this case, MPtrj), thus requiring only a modest dataset to achieve sufficient accuracy in specialized tasks. The combination of active learning strategies using cheap UIP trajectories to generate synthetic datasets for pre-training MLIPs can also greatly increase the efficiency of generating new datasets and expanding current ones [29, 30].
The performance of MLIPs for surfaces is important because, in materials science, a plethora of different phenomena take place on the surface of materials. Examples are adsorption and absorption of molecules and gases, catalysis, electrochemical processes such as water splitting, deposition and growth of new materials and phases such as heterostructures and nanoparticles. The generalization including both extensivity and transferability of MLIPs allows for simulations and understanding of ever larger and complex systems at an efficient computational cost. This enables the study of nanoscale and quantum materials systems such as 2D materials and their twisted interfaces, as well as more general global challenges involving molecules and materials, such as in energy, health, and the environment [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. As these models become more used, understanding their application limits and broadening them is of great importance to accelerate materials science research and applications of these models to different systems and use cases.
We envision that a promising step forward for the community is to create a universal training dataset for models to be trained upon, including not only all elements, but more importantly, intelligent sampling of the full combinatorial materials space composed of molecules, bulk solids, surfaces, interfaces, defects, multi-element alloys, and so on. Even though the currently available materials databases are a great starting point, they have been created with specific tasks in mind, focused on materials discovery. Compared to the huge size of the materials space, they still represent a very scarce fraction of all possible combinations. From a broader perspective, incorporating additional properties such as magnetic moments [18] or charge densities [41, 42], can be promising routes for further advancements [16]. There are still plenty of advances ahead.
IV Computational Details
IV.1 Datasets
The Materials Project surface dataset [24] was accessed on Feb. 2023 through Materials Project’s API [43]. We call $*{}^{*}$ the MP data that is filtered for the following set of elements from the MPtrj [44] dataset: Ni (96 structures), Cu (55 structures), Li (83 structures), Mo (25 structures), Si (559 structures), Ge (225 structures). Accordingly, for MTP and NequIP, a different model is fit for each element. The entire $*{}^{*}$ dataset is used for training and the test is performed on the MP surfaces dataset. The so-called JPCA dataset is the one made available from ref. [25].
IV.2 Training specialized potentials
IV.2.1 NequIP models
[…]
IV.2.2 MTP models
The moment tensor potential (MTP) models construct a contracted rotationally invariant representation of the atomic local environments with tensors from a set of basis functions, these are used to build a linear expansion of the potential energy as a function of the atomic representation [27]. The MTP models trained with the $*{}^{*}$ dataset are trained using the same radius cutoff as the models in ref. [25] for each element, while other hyperparameters are kept as the default from the maml package [46], while the so-called MTP [JPCA] models are the models made available from ref. [25]. Trained MTP models are available as Supporting Information.
IV.3 Fine-tuning MACE model
We fine-tune the MACE model with the aid of the training scripts provided with the MACE code package. The initial model for fine-tuning MACE is the MACE-MP-0 model trained on ref. [19], used throughout the text as MACE. The data used for fine-tuning is available from ref. [25]. Further details and command options, as well as the trained model, are made available as Supporting Information.
Supporting Information
Dataset exploration: surface energy data distribution; Evaluation of UIPs: correlation between UIPs’ predictions, understanding the errors by system chemistry; Comparison between UIPs, specialized MLIPs, and fine-tuned UIP: parity plot of all models. All trained models are available as files with instructions on their use in a compressed .zip file.
Acknowledgements
The authors acknowledge support from CNPq project no. 422069/2023-0 and CNPq - INCT (National Institute of Science and Technology on Materials Informatics), grants no. 371610/2023-0. LPMF is a CNPq scholarship holder - Brazil (150597/2023-1).